OpenStudy (bbb911):

This is Geometry: Find the surface area of a sphere whose radius is 7 millimeters. Round to the nearest hundredth.

OpenStudy (mathteacher1729):

Step 1 ) print this: http://mdk12.org/instruction/curriculum/hsa/geometry/math_reference_sheet.html

OpenStudy (anonymous):

The surface area of a sphere is given by: A=4(pi)(r^2), all we do is plug the radius 7 into the equation. (4)(pi)(25)=314.16

OpenStudy (bbb911):

here is the multiple choice for the q. A. 813.55 B. 615.75 C. 205.25 D. 175.93

OpenStudy (amistre64):

Sa = 4pi r^2 Sa = 4pi 7^2 Sa = 4pi 49 Sa = 196pi = abt. 615.75

OpenStudy (bbb911):

Thanks :D

OpenStudy (anonymous):

my bad i plugged in5 instead of 7

OpenStudy (bbb911):

its okay :)